An elevator starts from rest and moves upward , accelerating at a rate of 1.2m/s^2 until it reaches a speed of 7.8m/s, which it then maintains. Two seconds after the elevator begins to move, a man standing 12m above the initial position of the top of the elevator throws a ball upward with an initial velocity of 20m/s. Determine the time when the ball will hit the elevator.

Grade:10

2 Answers

Aman Bansal
11 years ago

Dear Pratiek

Try to solve the problem using relative velocity and newton equations

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

 

 

 

Neha Gaur
11 years ago

TIME TAKEN BY ball to reach highest point-> 0=20-9.8t =>t=2.04s

dist. covered by ball to reach the highest point->0=400-19.6s =>s=20.4m

dist covered by lift in (2+2.04)s=4.08m

distance between ball at 4.04sfrom start of lift=20.4+12-4.08=28.32

let after t seconds from falling of ball ,it hits elevator

then dist. covered by ball in t seconds=28.32m -dist covered by lift in t sec

                                          1/2gt^2=28.32m - (1/4)t^2

                                          (4.9 + 0.25)t^2=28.32

                                           5.15 t^2 =28.32

                                            t=2.34s

so ball hits elevator at (2+2.04+2.32)=6.36s